OpenStudy (anonymous):

0.7x+5le1.5x-2 please help

OpenStudy (anonymous):

\[0.7x \le1.5x-2\]

OpenStudy (anonymous):

2<=1.5x-0.7x 2<=0.8x 2.5<=x

OpenStudy (anonymous):

would it b \[\left\{ x \left| x <2.5 \right| \right\}\]

OpenStudy (anonymous):

not so sure><

OpenStudy (anonymous):

i dont know if it flips or not either

OpenStudy (anonymous):

\[2.5 \le x 0r x \ge 2.5\]

OpenStudy (anonymous):

so the sign does not flip?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

\[x \ge2.5\] The sign only flips if dividing by a negative number